OpenStudy (anonymous):

Consider the system {6q+3r=-1.5 q-5r=21.75 a) write the system as a matrix equation b) solve the system

OpenStudy (anonymous):

\[\begin{pmatrix} 6&3\\1&-5 \end{pmatrix}\begin{pmatrix}q\\r\end{pmatrix}=\begin{pmatrix}-1.5\\21.75\end{pmatrix}\]

OpenStudy (anonymous):

I got that part but what about b?

OpenStudy (anonymous):

You have an equation of the form \[{\bf A}{\bf x}={\bf b}\] where \(\bf A\) is the coefficient matrix (leftmost above), \(\bf x\) is the variable matrix, and \(\bf b\) is the constant matrix. What you can do is multiply both sides by the inverse of \(\bf A\): \[{\bf A}^{-1}{\bf A}{\bf x}={\bf A}^{-1}{\bf b}\] and you have \({\bf A}^{-1}{\bf A}={\bf I}\), where \(\bf I\) is the identity matrix, and you're left with \[{\bf x}={\bf A}^{-1}{\bf b}\] In the context of your problem, you have \[\begin{pmatrix} q\\r \end{pmatrix}=\begin{pmatrix} 6&3\\1&-5 \end{pmatrix}^{-1}\begin{pmatrix} -1.5\\21.75 \end{pmatrix}\] Do you know how to find the inverse of a 2x2 matrix?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

I believe you said you know how to find the inverse...

OpenStudy (anonymous):

okay for the first system?

OpenStudy (radar):

What did you get for the solution? q & r ??

OpenStudy (anonymous):

i dont know

OpenStudy (radar):

I assume they want you to solve it using the matrix method.

OpenStudy (radar):

 solve for r in a similar manner by substituting the constants for the coefficients of r